# SWE-Fixer: Training Open-Source LLMs for Effective and Efficient GitHub Issue Resolution

### v2

@@ Abstract
Large Language Models (LLMs) have demonstrated remarkable proficiency across a variety of complex tasks. One significant application of LLMs is in tackling software engineering challenges, particularly in resolving real-world tasks on GitHub by fixing code based on the issues reported by the users. However, many current approaches rely on proprietary LLMs, which limits reproducibility, accessibility, and transparency. The critical components of LLMs for addressing software engineering issues and how their capabilities can be effectively enhanced remain unclear. To address these challenges, we introduce SWE-Fixer, a novel open-source framework designed to effectively and efficiently resolve GitHub issues. SWE-Fixer comprises two essential modules: a code file retrieval module and a code editing module. The retrieval module employs BM25 along with a lightweight model to achieve coarse-to-fine file retrieval. Subsequently, the code editing module utilizes the other model to generate patches for the identified files. To mitigate the lack of publicly available datasets, we compile an extensive dataset that includes 110K GitHub issues along with their corresponding patches and train the two models of SWE-Fixer separately. We assess our approach on the SWE-Bench Lite and Verified benchmarks, achieving state-of-the-art performance among open-source models with scores of 24.7% and 32.8%, respectively. Additionally, our approach requires only two model calls per instance, making it significantly more efficient than existing methods. These results highlight the effectiveness of SWE-Fixer in real-world code-fixing scenarios. We will make our model, dataset, and code publicly available at https://github.com/InternLM/SWE-Fixer.
@@ 1 Introduction
Current approaches to resolving the Github issues in SWE-Bench can be broadly categorized into two main paradigms: agent (Wang et al., 2024b; Ma et al., 2024; Albert Örwall, 2024) and pipeline (Xia et al., 2024; AppMap, 2024). Agent-based systems rely on LLMs dynamically determining the next action, allowing them to autonomously explore a codebase and resolve issues. In contrast, pipeline-based methods guide LLMs through a series of predefined steps, such as initially identifying the defective files and subsequently editing them to resolve the issue. Despite significant advancements in agent and pipeline-based frameworks, most existing solutions depend on powerful proprietary models such as GPT-4o and Claude-3.5-Sonnet. While these solutions are effective, they are often costly and lack transparency, preventing a deeper understanding and further improvements in the problem-solving ability of LLMs. This motivates us to develop open-source LLMs to resolve the Github issues effectively and efficiently.
@@
SWE-Fixer achieves SOTA performance among open-source models. We evaluate SWE-Fixer on both SWE-Bench Lite and Verified, achieving 24.7% and 32.8% Best@1 performance, respectively, which are the highest among open-source model-based approaches. Furthermore, SWE-Fixer requires only two model calls per instance, making it significantly more efficient than existing open-source frameworks while maintaining superior performance. Notably, compared to frameworks built on proprietary models, SWE-Fixer outperforms several methods on both benchmarks, particularly those based on GPT-4, GPT-4o, and Claude-3-Opus, demonstrating exceptional efficiency and effectiveness. We also conduct a comprehensive study on training configurations for both tasks to provide insights into optimizing future systems. Additionally, the trained retriever and editor from SWE-Fixer have the potential to serve as modular components in agent-based systems, further enhancing the efficiency and effectiveness of existing agent-based approaches. Our contributions can be summarized as follows:
  • State-of-the-art performance: We propose a novel pipeline-based method that leverages open-source models, achieving state-of-the-art Best@1 performance on SWE-Bench Lite and Verified compared to other open-source model-based approaches.
@@ 3.3 Code Editing
To assist the model in generating valid code modifications, we define a structured output that includes three key components (see Appendix G). The first component is the file path, which specifies the location of the file to be modified. The second is the original code block, which includes the specific code snippet to be edited along with its line numbers. The third component is the modified code block, which represents the final modification result but excludes line numbers as the new line numbers are difficult for the model to calculate. The input includes complete file content with line numbers. This approach ensures that the model can efficiently identify the lines to edit and produce valid modifications. In contrast, the standard patch format [^1] https://git-scm.com/docs/diff-generate-patch requires line number calculations in the output hunk, which adds complexity. Our structured output simplifies training and can be transformed to patches automatically for evaluation.
@@ 4.2 Chain-of-Thought Data Construction
To address these challenges, we adopt a methodology inspired by the rationalization approach in Zelikman et al. (2022). Specifically, we use the gold patches as part of the input to guide the teacher model in generating both the reasoning process and the corresponding patches. The model is tasked with producing a reasoning chain and code patch as if it were unaware of the correct answer. Detailed prompts can be found in Appendix E. We use GPT-4o for the generation, and the resulting reasoning chains are generally coherent and sound.
Table 1: Performance of various models/methods on SWE-Bench Lite and Verified. : Finetuned open-source models. : All results are reported as Pass@1 or Best@1 in this table, except for SWE-Gym, which uses a specially trained 32B verifier.
@@
SWE-Fixer (Ours) SWE-Fixer-72B Pipeline 32.8 24.7
@@ 5.1 Experimental Setup
We finetune the Qwen2.5 (Qwen et al., 2024) base series models to implement SWE-Fixer, which consists of a 7B code retriever and a 72B code editor. The training is conducted on 96 Nvidia A800 GPUs using the xtuner-lite framework (Contributors, 2023), with a global batch size of 96 and a 64K context window. We evaluate SWE-Fixer on two datasets: SWE-Bench Lite and SWE-Bench Verified. SWE-Bench Lite, part of the official SWE-Bench benchmark (Jimenez et al., 2023), consists of 300 selected instances optimized for efficient evaluation. SWE-Bench Verified, proposed by OpenAI, is a human-validated subset that offers a more reliable evaluation. Each instance includes a real-world GitHub issue description paired with its corresponding codebase. Model-generated code patches are assessed using developer-written unit tests, and accuracy is calculated as the percentage of instances successfully resolved.
@@ 5.2 Dataset Preparation
Table 2: Performance comparison of different methods with open-source models on SWE-Bench Verified and Lite. ‘#Model Calls’ represents the number of model calls required per instance. : The minimum number of model calls needed. : The average number of model calls needed. : See Appendix C for the detailed calculation.
@@
SWE-Fixer (Ours) SWE-Fixer-72B 32.8 24.7 2
@@ 5.3 Main Results
Among the approaches based on open-source models, our SWE-Fixer framework achieves the highest Best@1 performance. While SWE-Gym-32B (Pan et al., 2024) attains a higher Best@8 score on SWE-Bench Lite with its specially trained 32B verifier, it still falls behind SWE-Fixer on SWE-Bench Verified. SWE-Gym represents a promising direction by leveraging reinforcement learning in executable environments. However, its heavy reliance on human effort for environment construction limits its scalability. SWE-Fixer outperforms SWE-SynInfer, an agent-based method using fine-tuned models, on both SWE-Bench Lite and Verified.
Beyond effectiveness, our method also demonstrates exceptional efficiency. As shown in Table 2, apart from the RAG method, which performs poorly, SWE-Fixer requires only two model calls (7B retrieval + 72B editing) per instance, making it significantly more efficient than other methods while maintaining state-of-the-art performance. Notably, SWE-Search requires at least 200 model calls per instance yet achieves the same score as SWE-Fixer on SWE-Bench Lite, making it prohibitively expensive in comparison.
@@ 5.4.2 Code Editing
Enhanced location information improves performance. Including line numbers acts as an anchor, helping LLMs locate and edit specific code snippets more effectively, which improves performance. Additional experiments on more fine-grained settings can be found in Appendix D.
@@ Appendix A Training Data Collection
Appendix B Post-processing
Appendix C Model Calls Calculation for SWE-Seach
Appendix D Ablation on More Fine-Grained Retrieval Strategies
@@
Appendix E CoT Generation Prompt
@@
Appendix F File Documentation
@@
Appendix G Edit Output Patch Example

### v3

@@ Abstract
Large Language Models (LLMs) have demonstrated remarkable proficiency across a variety of complex tasks. One significant application of LLMs is in tackling software engineering challenges, particularly in resolving real-world tasks on GitHub by fixing code based on the issues reported by the users. However, many current approaches rely on proprietary LLMs, which limits reproducibility, accessibility, and transparency. The critical components of LLMs for addressing software engineering issues and how their capabilities can be effectively enhanced remain unclear. To address these challenges, we introduce SWE-Fixer, a novel open-source framework designed to effectively and efficiently resolve GitHub issues. SWE-Fixer comprises two essential modules: a code file retrieval module and a code editing module. The retrieval module employs BM25 along with a lightweight model to achieve coarse-to-fine file retrieval. Subsequently, the code editing module utilizes the other model to generate patches for the identified files. To mitigate the lack of publicly available datasets, we compile an extensive dataset that includes 110K GitHub issues along with their corresponding patches and train the two models of SWE-Fixer separately. We assess our approach on the SWE-Bench Lite and Verified benchmarks, achieving competitive performance among open-source models with scores of 22.0% and 30.2%. Furthermore, SWE-Fixer reaches state-of-the-art performance (24.7% on Lite and 32.8% on Verified) with PASS_TO_PASS (P2P) filtering. Additionally, our approach requires only two model calls per instance, making it significantly more efficient than existing methods. These results highlight the effectiveness of SWE-Fixer in real-world code-fixing scenarios. We will make our model, dataset, and code publicly available at https://github.com/InternLM/SWE-Fixer.
@@ 1 Introduction
Current approaches to resolving the Github issues in SWE-Bench can be broadly categorized into two main paradigms: agent (Wang et al., 2024b; Ma et al., 2024; Albert Örwall, 2024) and pipeline (Xia et al., 2024; AppMap, 2024; Li et al., 2025). Agent-based systems rely on LLMs dynamically determining the next action, allowing them to autonomously explore a codebase and resolve issues. In contrast, pipeline-based methods guide LLMs through a series of predefined steps, such as initially identifying the defective files and subsequently editing them to resolve the issue. Despite significant advancements in agent and pipeline-based frameworks, most existing solutions depend on powerful proprietary models such as GPT-4o and Claude-3.5-Sonnet. While these solutions are effective, they are often costly and lack transparency, preventing a deeper understanding and further improvements in the problem-solving ability of LLMs. This motivates us to develop open-source LLMs to resolve the Github issues effectively and efficiently.
@@
SWE-Fixer achieves 22.0% on SWE-Bench Lite and 30.2% on SWE-Bench Verified, performing comparably to existing SOTA open-source models. Additionally, with P2P filtering, SWE-Fixer achieves Best@1 performance of 24.7% on SWE-Bench Lite and 32.8% on SWE-Bench Verified. These results set a new SOTA for open-source model-based approaches. Furthermore, SWE-Fixer requires only two model calls per instance, making it significantly more efficient than existing open-source frameworks while maintaining superior performance. Notably, compared to frameworks built on proprietary models, SWE-Fixer outperforms several methods on both benchmarks, particularly those based on GPT-4, GPT-4o, and Claude-3-Opus, demonstrating exceptional efficiency and effectiveness. We also conduct a comprehensive study on training configurations for both tasks to provide insights into optimizing future systems. Additionally, the trained retriever and editor from SWE-Fixer have the potential to serve as modular components in agent-based systems, further enhancing the efficiency and effectiveness of existing agent-based approaches. Our contributions can be summarized as follows:
  • State-of-the-art performance: We propose a novel pipeline-based method that leverages open-source models, achieving state-of-the-art Best@1 performance on SWE-Bench Lite and Verified with P2P filtering compared to other open-source model-based approaches.
@@ 3.3 Code Editing
To assist the model in generating valid code modifications, we define a structured output that includes three key components (see Appendix H). The first component is the file path, which specifies the location of the file to be modified. The second is the original code block, which includes the specific code snippet to be edited along with its line numbers. The third component is the modified code block, which represents the final modification result but excludes line numbers as the new line numbers are difficult for the model to calculate. The input includes complete file content with line numbers. This approach ensures that the model can efficiently identify the lines to edit and produce valid modifications. In contrast, the standard patch format [^1] https://git-scm.com/docs/diff-generate-patch requires line number calculations in the output hunk, which adds complexity. Our structured output simplifies training and can be transformed to patches automatically for evaluation.
@@ 4.2 Chain-of-Thought Data Construction
To address these challenges, we adopt a methodology inspired by the rationalization approach in Zelikman et al. (2022). Specifically, we use the gold patches as part of the input to guide the teacher model in generating both the reasoning process and the corresponding patches. The model is tasked with producing a reasoning chain and code patch as if it were unaware of the correct answer. Detailed prompts can be found in Appendix F. We use GPT-4o for the generation, and the resulting reasoning chains are generally coherent and sound.
Table 1: Performance of various models/methods on SWE-Bench Lite and Verified. : Finetuned open-source models. : All results are reported as Pass@1 or Best@1 in this table, except for SWE-Gym, which uses a specially trained 32B verifier. P2P filtering means that we use the P2P tests in the repository to discard patches that break them, retaining only those that pass all P2P tests.
@@
SWE-Fixer SWE-Fixer-72B Pipeline 30.2 22.0
SWE-Fixer + P2P Filtering SWE-Fixer-72B Pipeline 32.8 24.7
@@ 5.1 Experimental Setup
We finetune the Qwen2.5 (Qwen et al., 2024) base series models to implement SWE-Fixer, which consists of a 7B code retriever and a 72B code editor. The training is conducted on 96 Nvidia A800 GPUs using the xtuner-lite framework (Contributors, 2023), with a global batch size of 96 and a 64K context window. We evaluate SWE-Fixer on two datasets: SWE-Bench Lite and SWE-Bench Verified. SWE-Bench Lite, part of the official SWE-Bench benchmark (Jimenez et al., 2023), consists of 300 selected instances optimized for efficient evaluation. SWE-Bench Verified, proposed by OpenAI, is a human-validated subset that offers a more reliable evaluation. Each instance includes a real-world GitHub issue description paired with its corresponding codebase. Model-generated code patches are assessed using developer-written unit tests, and accuracy is calculated as the percentage of instances successfully resolved. Some instances include PASS_TO_PASS (P2P) tests, which can be used to filter patches and ensure they do not affect unrelated functionality in the repository. This filtering is optional, and further details are provided in Appendix C.
@@ 5.2 Dataset Preparation
Table 2: Performance comparison of different methods with open-source models on SWE-Bench Verified and Lite. ‘#Model Calls’ represents the number of model calls required per instance. : The minimum number of model calls needed. : The average number of model calls needed. : See Appendix D for the detailed calculation.
@@
SWE-Fixer SWE-Fixer-72B 30.2 22.0 2
SWE-Fixer + P2P Filtering SWE-Fixer-72B 32.8 24.7 2
@@ 5.3 Main Results
Our SWE-Fixer framework achieves comparable performance among the approaches based on open-source models. While SWE-Gym-32B (Pan et al., 2024) attains a higher Best@8 score on SWE-Bench Lite with its specially trained 32B verifier, it still falls behind SWE-Fixer on SWE-Bench Verified. SWE-Gym represents a promising direction by leveraging reinforcement learning in executable environments. However, its heavy reliance on human effort for environment construction limits its scalability. SWE-Fixer achieves the same performance as SWE-Synifer, an agent-based approach that requires significantly more model calls and is much more complex. With P2P tests filtering, SWE-Fixer achieves the highest Best@1 performance among all open-source model-based methods.
Beyond effectiveness, our method also demonstrates exceptional efficiency. As shown in Table 2, apart from the RAG method, which performs poorly, SWE-Fixer requires only two model calls (7B retrieval + 72B editing) per instance, making it significantly more efficient than other methods while maintaining comparable performance. Notably, SWE-Search requires at least 200 model calls per instance yet achieves only a slightly higher score than SWE-Fixer on SWE-Bench Lite without P2P filtering, rendering it prohibitively expensive by comparison.
@@ 5.4.2 Code Editing
Enhanced location information improves performance. Including line numbers acts as an anchor, helping LLMs locate and edit specific code snippets more effectively, which improves performance. Additional experiments on more fine-grained settings can be found in Appendix E.
@@ Appendix A Training Data Collection
Appendix B Post-processing
Appendix C P2P Filtering
Pass-to-Pass (P2P) filtering refers to using regression tests from each code repository to validate the correctness of generated patches. Specifically, a P2P test is a test case that passes both before and after applying the gold patch, and can be considered a check for the correctness of unrelated functionality in the repository. SWE-Bench provides a list of such tests for each issue. Following the setup of earlier works such as Agentless v1 Xia et al. (2024) and Aider, in some experiments, we apply P2P filtering in our inference stage. Concretely, we apply the generated patch to the repository and verify whether it breaks any P2P tests. If the patch passes all P2P tests, it is retained; otherwise, we resample a new patch. We first use greedy sampling to generate the patch. If the patch fails, a temperature of 0.7 is applied to generate more creative outputs. On instances where the patch ultimately passes, the average number of generation attempts is 1.15 on SWE-Bench Lite and 6.73 on SWE-Bench Verified. Given the ongoing community discussion on the SWE-Bench GitHub repository regarding whether P2P filtering should be used during inference, we report both P2P-filtered and non-filtered results in our paper for transparency and comparison.
Appendix D Model Calls Calculation for SWE-Seach
Appendix E Ablation on More Fine-Grained Retrieval Strategies
@@
Appendix F CoT Generation Prompt
@@
Appendix G File Documentation
@@
Appendix H Edit Output Patch Example